### v1

# A New Era in Computational Pathology: A Survey on Foundation and Vision-Language Models

Abstract
Recent advances in deep learning have completely transformed the domain of computational pathology (CPath), which in turn altered the diagnostic workflow of pathologists by integrating foundation models (FMs) and vision-language models (VLMs) in their assessment and decision-making process. FMs overcome the limitations of existing deep learning approaches in CPath by learning a representation space that can be adapted to a wide variety of downstream tasks without explicit supervision. VLMs allow pathology reports written in natural language to be used as a rich semantic information source to improve existing models as well as generate predictions in natural language form. In this survey, a holistic and systematic overview of recent innovations in FMs and VLMs in CPath is presented. Furthermore, the tools, datasets and training schemes for these models are summarized in addition to categorizing them into distinct groups. This extensive survey highlights the current trends in CPath and the way it is going to be transformed through FMs and VLMs in the future.
Index Terms:
Computational Pathology, Foundation Models, Multi-Modal, Vision-Language Models
I Introduction
In recent years there has been a surge of artificial intelligence (AI) based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] in CPath owing to wide adoption of digital slide scanners. As a result, large-scale curation and annotation [19, 20] of whole slide images (WSIs) has been made possible which ensured adequate data to train these AI-based models. The goal of these AI-based models is to automate and expedite the diagnosis and prognosis process of CPath. The traditional diagnosis process in CPath is time-consuming and requires experts with extensive domain knowledge. In addition, the wide variation of pathology and heterogeneity between tasks makes it difficult to come up with a unified general approach.
Several research studies have addressed this lack of a unified approach and among the proposed methods, the FMs have gained a lot of attention in recent years [21, 22, 23, 24, 25, 26]. FMs leverage self-supervised learning (SSL) [27] schemes to learn a rich representation in a task-agnostic manner. Owing to self-supervised pre-training (SSPT), FMs do not require large-scale annotated data which is hard to come by in CPath. Furthermore, these models can be trained with a diverse selection of datasets containing tissue samples from different organs and associated with different cancer types, scanner types, etc. As a result, the resultant pre-trained model can easily be utilized in a wide range of downstream tasks while maintaining robustness to extreme variation in tissue samples.
Figure 1: Number of publications in FMs and VLMs in pathology (from Google Scholar). The search was done with the keywords “vision-language” + “pathology” for VLMs statistics and “foundation models”+ “pathology” for FMs statistics.
Figure 2: Outline of the major challenges in CPath (challenges in data collection, challenges in data annotation, lack of diverse data, large number of tasks and challenges in deep learning architectures). Several causes and consequences for each challenge are outlined in addition to how FMs and VLMs address these challenges.
The impact of FMs in CPath can be amplified by integrating the power of VLMs [28] which surged in popularity after the introduction of contrastive language-image pre-training (CLIP) [29] model by OpenAI. Pathology reports, books, education videos, etc are a rich source of semantic information that can be utilized by VLMs to significantly boost performance which is not possible with vision-only models. When used in conjunction with FMs they can perform like AI pathologists capable of performing a vast array of tasks as evidenced by recent research works [25]. Even though CPath is a specialized field, there is a large increase in the number of publications focusing on FMs and VLMs as shown in Fig. 1 which indicates the future direction of this field.
To appreciate the full impact of FMs and VLMs in CPath, the major challenges in CPath are outlined below. In Fig. 2 key points of these challenges are summarized. All these challenges are addressed in some form by FMs and VLMs as mentioned in Fig. 2.
I-A Scope of the Review
In this review, the main emphasis is put on the application of FMs and VLMs in CPath, especially the details of their architectures and training schemes. Note that these two categories are not mutually exclusive, meaning that some research articles belong to both categories which are vision-language foundation models (VLFMs). In addition, details of multi-modal datasets are summarized with a focus on vision and language as the modalities.
TABLE I: Number of Surveyed articles published in top journals and conferences within 2023-2024
Journal/Conference Venue Counts
Nature/Nature Medicine 6
CVPR 8
NeurIPS 2
ECCV 2
MICCAI/MICCAI Workshop 4
AAAI 1
Elsevier/Springer 3
Several self-imposed rules and restrictions were used as guidelines throughout the review to ensure the scope of the paper is maintained.
  1. 1.
    First, articles that focus solely on pathology are included and articles that focus on other areas of the biomedical domain are excluded. As an example, articles like BiomedCLIP [30] with pathology as a subsection of the research are not included in the review.
  2. 2.
    Secondly, only vision-language models are included and articles that use other modalities along with vision are excluded. As an example, along with vision, transcriptomics[31] can be used to solve pathological tasks. However, such papers are excluded to maintain the scope.
  3. 3.
    Both peer-reviewed articles and pre-print articles are included in the survey. Among the peer-reviewed articles, a significant number of articles are published in top-tier journals and conferences as shown in Table I.
Figure 3: Visualization of the timeline of recently published work in CPath utilizing FMs and VLMs as well as multi-modal datasets. To maintain transparency we clearly annotate research articles that have been peer-reviewed and articles that are available as pre-print. Furthermore, high-impact pioneering research works published in prominent journals are highlighted. For pre-prints if there are multiple version, the latest version and the corresponding date is used.
I-B Contribution and Organization
In the past few years, quite a few review articles [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] have been published focusing on computational pathology. Most of these reviews include articles in the digital and computational pathology domain and the application of deep learning as a whole rather than focusing on a specific subtopic. There are a handful of papers that focus on specific architecture like MIL [32], graph-based models [33, 17], transformer-based models [13, 10], LLMs [18], etc. The contribution of this review article is mentioned below:
  1. 1.
    To the best of our knowledge, this is the first review article to summarize recent advances (most articles from 2023-2024) in foundation and vision-language models (Section III and IV). The closest peer-reviewed survey article [8] mostly provides a high-level overview without going into the details of foundation models.
  2. 2.
    An exhaustive list of multi-modal datasets (Section II) in computational pathology that are being used or can be used in vision-language research is outlined.
  3. 3.
    Given the diverse datasets and architectures, descriptions for individual research are provided in tabular format (Table II, Table IV, Table V, Table VIII) so it is easier for readers to follow.
  4. 4.
    A categorized and annotated timeline of the survey articles (Fig. 3) which provides a clear idea of the evolution of the FMs and VLMs in CPath.
The rest of the article is organized as follows. In section II existing multi-modal datasets are listed along with details about their source, pre-processing techniques, etc. In section III existing FMs in CPath are outlined and descriptions for the vision and vision-language pre-training schemes for these FMs are provided. In section IV an extensive list of VLMs in CPath is provided along with details about their architecture, utilized datasets and contribution. Finally, in section V the paper is concluded.
II Multi-Modal Datasets in Pathology
In this section, a comprehensive summary of the existing multi-modal datasets available in CPath is provided. The modalities taken into account are vision and language. The one article that has additional modalities is CR-PathNarratives which is included because it was referenced in many other VLMs articles. All key information regarding each dataset is summarized in Table II.
Figure 4: Different components of multi-modal datasets in computational pathology: Type of datasets, source of data, annotation and pre-processing
To have a comprehensive understanding of the existing multi-modal datasets for pathology, three different components need to be considered as shown in Fig. 4.
TABLE II: Summary of Multi-Modal Datasets in Pathology
Dataset Type of Data Size Image and Text Source Method of Generation/Procurement Other Utilized Dataset/Models for Generation/Procurement Dataset Used By Variation/ Subset/ Extension Availability (Linked)
PathText [34] WSI and question-text pairs 1,041 WSI 9,009 WSI-text pairs WSI and pathology report from TCGA-BRCA Prompting LLMs, OCR, manual annotation, classifier MI-Gen [34]
WSI-VQA [35] WSI and Q/A pairs 977 WSIs 8,672 question-answer pairs (slide-level) with average 8.9 Q/A pairs per WSI WSI and pathology report from TCGA-BRCA Prompting LLMs and template matching heuristics Model/Framework: GPT-4 [36] Wsi2Text Transformer (W2T) [35] 4535 close-ended VQA subset and 4137 open-ended VQA subset
PathGen-1.6M [37] Image-caption pairs 1.6 million WSI and pathology report from TCGA Multi-agent collaboration and caption generation with LMMs Dataset: PathCap [38], Quilt-1M [39], OpenPath [40] Model/Framework: LLaVA-v1.5 [41], Vicuna [42] OpenCLIP [43], GPT-4 [36] PathGen-CLIP[37], PathGen-LLVA [37] 200K instruction-tuning data (Extension)
PathMMU [44] Image and Q/A pairs 24,067 pathology images 33,428 Q/A PubMed, pathology atlas, educational videos, pathologist-shared images with explanations on Twitter Prompting GPT-4, Heuristics, Annotation by seven pathologists Dataset: Quilt-1M [39], OpenPath [40] Model/Framework: GPT-4 [36], YOLOv6 [45] PubMed, SocialPath EduContent, Atlas, PathCLS (Subsets)
Quilt-Instruct [46] Instruction-tuning question-answer pairs 107,131 histopathology-specific Q/A pairs Over 1,000 hours of 4,149 educational histopathology videos from YouTube Prompting GPT-4, hand-crafted algorithms for extraction of video frames and spatial annotation Dataset: Quilt-1M [39] Model/Framework: GPT-4 [36] QUILT-LLAVA[46] Quilt-VQA with 985 images 1,283 Q/A pairs where 940 are open-set and 343 closed-set
QUILT [39] Image-text pairs 437,878 images aligned with 802,144 text pairs 1,087 hours of 4,504 narrative educational histopathology videos from YouTube Prompting LLMs, hand-crafted algorithms, human knowledge databases (UMLS), automatic speech recognition Dataset: OpenPath [40], PubMed, LAION-5B [47] (For Quit-1M) Model/Framework: Whisper [48], GPT-3.5, inaSpeechSegmenter [49], langdetect [50] QUILT-Net[39] Quilt-1M with 1 million image-text pair with additional data from LAION, Twitter, and PubMed (Extension)
PathQABench [25] ROI-annotated WSI and Q/A 48 H&E WSIs (25 WSIs from PathQABench-Private 23 WSIs from PathQABench-Public) + 48 close-ended Q/A 115 open-ended Q/A PathQABench-Private from private in-house cases PathQABench-Public from TCGA cases Expert-pathologists curated and annotated PathChat[25] (Subsets) PathQABench-Public and PathQABench-Private PathChatInstruct 456,916 instruction-tuning dataset (Extension)
PathCap [38] Image-caption pairs 207K PubMed, internal pathology guidelines books, annotation by expert cytologists Parsing from PubMed, image processing with YOLOv7, ConvNeXt, PLIP, caption refinement and text processing with ChatGPT Dataset: PubMed Model/Framework: YOLOv7 [51], ConvNeXt [52] ChatGPT [53], PLIP [40] PathAsst [38] PathInstruct 180K pathology multimodal instruction-following samples (Extension)
OpenPath [40] Image-text pairs 208,414 image–text pairs 116,504 image–text pairs from Twitter posts, 59,869 image–text pairs from replies, 32,041 image–text pairs from LAION-5B Pathology image classifier to exclude non-pathology images, CLIP image embeddings with cosine similarity to create PathLAION, other hand-crafted heuristics Dataset: LAION-5B [47] Model/Framework: CLIP [29], langdetect [50] PLIP [40] PathLAION 32,041 pathology images from the LAION-5B dataset (Subset)
CR-PathNarratives [54] WSIs with multi-modal annotations 174 annotated colorectal WSIs Proprietary data from a hospital Annotation by eight pathologists which contains ROI, voice information and behavioral trajectory information of the annotators Model/Framework: PathNarrative annotation tool [54]
ARCH [55] Image-caption pairs 11,816 image-caption pairs PubMed and pathology textbooks Hand-crafted algorithms with tools like Pubmed Parser, Pdffigures 2.0 Model/Framework: Pubmed Parser [56] Pdffigures 2.0 [57]
PathVQA [58] Image and Q/A pairs 4,998 images 32,799 Q/A pairs Pathology textbooks, Pathology Education Informational Resource (PEIR) digital library Hand-crafted algorithms with tools like PyPDF2, PDFMiner, Stanford CoreNLP toolkit Model/Framework: PyPDF2, PDFMiner Stanford CoreNLP toolkit [59] 16,465 open-ended Q/A subset, 16,334 close-ended Q/A subset
Prov-Path [26] WSI and reports pairs 17,383 WSI-reports pairs Proprietary dataset from Providence Health System (PHS) K-means to generate four representative reports which are used to prompt GPT 3.5 to clean rest of the reports Model/Framework: GPT-3.5 Prov-GigaPath [26]
Dataset from CONCH  [24] Image-caption pairs 1,786,362 image-caption pairs PubMed, publicly available research articles, internal data from Mass General Brigham institution Hand-crafted workflow with YOLOv5, CLIP, BioGPT Model/Framework: YOLOv5 [60], CLIP BioGPT [61] CONCH [24] PMC-Path (data from PubMed) EDU (data extracted from educational notes)
Dataset from MI-Zero  [62] Image-caption pairs 33,480 image-caption pairs Publicly available educational resources combined with ARCH Hand-crafted algorithms to filter out non-pathology images Dataset: ARCH MI-Zero [62]
Availability: GitHub repositories (except for ARCH [55] and PathVQA [58]; direct dataset source is linked for these two papers) with the associated paper are linked if they are open-sourced and accessible. Some repositories only provide data generation and processing code (if the data is proprietary or requires an API) and some provide direct data sources through platforms such as Hugging Face Hub, Zenodo. Other Utilized Dataset/Models: Dataset/Models/Framework used only in the data generation/ processing part is mentioned. Downstream task dataset is not mentioned for this reason.
II-A Type of Datasets:
The first component is the type of data which can be broadly classified into $55$ categories. CR-PathNarratives [54] and Prov-Path [26] fall into the last category which is dissimilar to the rest of the datasets. In CR-PathNarratives, WSIs are associated with special annotated additional modalities. The additional modalities are voice and behavioral trajectory information of the pathologists who annotated the dataset. These additional modalities were collected by using the PathNarrative multimodal interactive annotation tool proposed in the same research work. The dataset Prov-Path utilizes WSI and the corresponding reports along with histopathology findings, cancer staging, genomic mutation profiles, etc collected by Providence Health System (PHS).
Figure 5: Comparison between the size of different multi-modal datasets. The size of the bubbles indicates the size of the data set and the number of elements in a data set is mentioned on the right of the bubbles. Color is used to indicate the different types.
The rest of the datasets can be categorized into the remaining four categories. The image-caption/ image-text pair category (PathGen-1.5M [37], Quilt-1M [39], OpenPath [40], PathCap [38], ARCH [55]) involves a low-to-medium quality image and an associated piece of text for that image. This text can be a short caption with a description of the image or a more elaborate description. ARCH is the earliest dataset in this category that utilized Pubmed and pathology textbooks to extract the texts. PathGen-1.5 M is the latest and largest dataset in this category, but unlike other datasets in this category, the images are patches extracted from WSIs. This is in contrast to other datasets in this category like OpenPath which is constructed with Twitter posts and replies and Quilt-1M which is constructed with frames extracted from educational pathology videos. However, the curation and annotation process of this category is much easier as it can automated with hand-crafted algorithms and heuristics. But it comes with the trade-off of noisy data as due to the automated process a lot of artifacts can be present in the data.
The WSI VQA/text category (PathText [34],WSI-VQA [35], PathQABench [25]) contains question-and-answer pairs or texts associated with WSIs. Comparatively, the generation process of this category is more difficult and often involves prompting LLMs to extract information or format information according to a certain template. The most common data source is the cancer genome atlas (TCGA) [19] which contains a large repository of WSI and patient report pairs. The VQA part can be of two types, close-ended question-answer pair and open-ended question-answer pair. Close-ended question-answer pairs are of a multiple-choice type or short-answer type. On the other hand, open-ended question-answer pairs contain answers that are in natural language form. Among the datasets in this category, PathQABench is unique as it contains ROI annotation of WSIs performed by expert pathologists. It has two subsets PathQABench-Public and PathQABench-Private. The former is publicly available as it was constructed with TCGA WSI and reports, and the latter was constructed with in-house data.
The next category, which is VQA (PathMMU [44], PathVQA [58], Quilt-VQA [46]) is similar to the previous category as it also contains close-ended and open-ended question-answer pairs, but the associated images are not WSIs but rather low-to-medium-quality images. Among these datasets, PathVQA is the first research to curate a pathology-specific VQA dataset. PathMMU is the latest and largest dataset in this category and it also provides explainability annotations with each answer. PathMMU utilized two previous datasets from the image-caption/image-text category OpenPath and Quilt-1M to generate a subset of the question-answer pair with images. Another category, which is the instruction-tuning dataset (Quilt-Instruct [46], PathInstruct [37], PathChatInstruct [25], extension of PathGen-1.6M [37] ) is unique kind of dataset, as this type of dataset is used to provide conversational ability to an existing multimodal model. The common workflow is that the instruction-tuning data set is applied in the last phase to fine-tune an already trained VLM. All of these datasets were created following the strategy mentioned in LLaVA [41] or LLaVa-1.5 [63]. In Fig. 5 a visual comparison of the size of each dataset in each category is provided.
II-B Source of Data:
The second component to consider is the source of data which largely dictates the third component, the annotation and pre-processing. PubMed is a common data source containing pathology images and captions/text. However, the quality of the data is not as high as that of the TCGA repository that contains WSIs and corresponding pathological reports. Other high-quality data sources that contain WSIs and pathology reports are in-house proprietary datasets. A unique data source, OpenPath, contains pathology images and texts from Twitter posts and replies associated with a large pathology community.
Figure 6: Subsets of the Quilt-1M dataset and the corresponding number of image-text pair for each subset.
Figure 7: Subsets of the PathMMU dataset and proportion for each subset. The EduContent and SocialPath subset is sourced from Quilt-1M and OpenPath dataset.
This data set was supplemented by pathology-specific data from the large-scale artificial intelligence open network (LAION) data repository. Pathology textbooks and atlas are also large knowledge sources that can be used to extract image caption/text pairs. In a couple of recent research [39, 46], educational histopathology videos on YouTube are being used as the source of pathology image and text pair. However, curation of this kind of dataset requires a series of hand-crafted algorithms and many external tools.
Based on the above discussion, it is apparent that there is a trade-off between the quality and volume of the data. Sources like PathQABench are superior in terms of quality, as they were curated by an expert pathologist. However as it requires explicit manual annotation and supervision, it is much harder to scale and hence small in terms of size.
Another key point is that most of these datasets contain other datasets as one of the subsets. An example of that is Quilt-1M which contains OpenPath, PubMed and LAION as subsets in addition to the proposed Quilt dataset which is the original proposed dataset. A visualization of Quilt-1M and its subsets is provided in Fig. 6. Another such example is the PathMMU dataset (shown in Fig. 7) which contains $55$ different subsets, PubMed, SocialPath, EduContent, Atlas and PathCLS each containing data from different sources.
II-C Annotation and Pre-Processing:
The third component is the annotation and pre-processing steps of data curation and generation. Among the surveyed articles, all employ a series of steps depending on the type of dataset and if the entire annotation and pre-processing pipeline is compared every article is unique. However, some specific steps are similar if the data source is the same. For example, all articles that utilize PubMed as a data source use some kind of parsing process [56] to parse and extract figures and texts. In addition, they employ light-weight classifiers and object detection architectures (YOLOv5 [60], YOLOv6 [45], YOLOv7 [51]) to distinguish between pathology and non-pathology images, detect and separate subfigures, etc. Another common approach is to prompt LLMs to format and refine captions/text or structure extracted information according to a pre-defined template. These LLMs include generalized LLMs like GPT-4, GPT-3.5, ChatGPT and also specialized LLMs like BioGPT. Another widely used strategy is using a trained CLIP-based model and using cosine similarity as a metric to classify pathology and non-pathology images.
Apart from the approaches mentioned above, there are a lot of other hand-crafted algorithms, heuristics and tools which are summarized in Table II.
III Foundation Model
Figure 8: A high-level visualization of different factors of variability in terms of organs, stain types, scanner types, magnification levels, preservation methods, downstream tasks, task levels, pre-training tissue sample sources, etc in foundation models. The research that puts the most emphasis on ensuring variability are RudolfV [64] and PLUTO [65]. Among the tissue sample sources for pre-training TCGA and PAIP are publicly accessible and the rest are proprietary.
In this section, an overview of existing FMs in CPath is provided. First, the characteristics of FMs are provided to remove any ambiguity for the later sections. Next, pre-training workflow and typical pre-training schemes are mentioned.
III-A Characteristics of FMs
A model can be classified as FM if it holds the following characteristics :
  1. 1.
    The first characteristic that is common to all FMs is the SSPT. The data used in the pre-training phase do not have any explicit label or annotation. Specifics of the SSL strategies are mentioned in section III-B.
  2. 2.
    The training goal of FMs is not to solve any specific task but rather to learn a general and rich representation space. For VFMs, it is a vision representation space and for VLFMs it is a vision-language representation space. The training of FMs is termed as “pre-training” as in later stages further training is required to optimize for a specific task.
  3. 3.
    In CPath, FMs are trained using large and diverse datasets that encompass tissue samples from different organs and anatomic sites. In addition, some research [64, 65] put effort into capturing diversity in terms of scanner types, magnification levels, stain types, preservation methods, etc. The idea is to capture the representation of a wide range of tissue and disease types and also making sure in downstream tasks are robust. In Fig. 8 a high-level visualization is provided highlighting different aspects of FMs.
  4. 4.
    Another characteristic is the size of models that typically have parameters on the scale of millions. A huge amount of computing resources is put into training these models involving multiple GPUs.
As shown in Fig. 3, the FMs are sectioned into three separate categories. One category encompasses VFMs [22] in CPath, the second category encompasses VLFMs in CPath and the last category utilizes these FMs by providing a benchmark [66, 67, 68, 69, 70], framework [71] or adapting FMs [72, 73, 74, 75, 76, 77]. Before going into the details of each category in the following section, an overview of the pre-training strategies of FMs is provided.
III-B Pre-training Workflow and Strategies
The typical pre-training workflow of FMs is shown in Fig. 9 which provides a high-level visualization of different phases in the workflow.
In this single diagram, both pre-training strategy for VFM and VLFM is shown. For VFMs, a vision module with some initial weight is leveraged in pre-training. The term “vision module” is used to generally represent a wide range of vision architectures and also modified versions of these architectures with additional layers. The most common architecture is variants of vision image transformers (ViTs) [78] which includes ViT-S, ViT-B, ViT-L and ViT-H. Though there are specialized architectures proposed in BEPH [79] and Prov-GigaPath [26] which uses BEiTv2 and GigaPath architecture respectively. Note that, some models initialize the architectures with ImageNet [80] weights to get an initial vision representation space which can be transformed through pre-training. In Table III, a summary of pre-training strategies in different phases is outlined.
Figure 9: Typical pre-training workflow in FMs. In general it starts with a vision module and a language module with randomly initialized weights or pre-set weights. VFMs only go through unimodal vision pre-training to learn a vision representation space. On the other hand, VLFMs can optionally go through unimodal pre-training for their vision and language modules.
TABLE III: Different pre-training phases and Corresponding SSPT Strategies
Pre-training phase Strategy
Unimodal vision pre-training Self-Distillation DINO [81], DINOv2 [82]
Contrastive Learning MoCov3 [83]
Masked Image Modeling (MIM) MAE [84], iBOT (MIM+Self distillation) [85]
Unimodal language pre-training No Specific Strategy
Vision-language pre-training CoCa [86], CLIP [29]
TABLE IV: Summary of Vision Pre-Training Strategy and Vision Pre-Training Dataset of Foundation Models in Pathology
Model Vision Pre-Training Strategy Vision Pre-Training Dataset Type of Foundation Model Availability (Linked)
Category Approach Architecture Source Size Stain Additional Information
GPFM [22] MIM + Self-Distillation+ (Expert Knowledge Distillation) Custom Custom architecture that uses other FMs including UNI, Phikon, CONCH $3333$ public datasets including TCGA, GTExPortal PAIP, CPTAC, etc Slides: 86,104 Tiles: 190,000,000 47 data sources 34 major tissue types Combines Knowledge of VFMs and VLFMs
Virchow 2 [21] Virchow 2G Self-Distillation DINOv2 ViT-H ViT-G Memorial Sloan Kettering Cancer Center (MSKCC) + Institutions worldwide Slides: 3,134,922 Tiles: H&E IHC 225,401 patients 493,332 cases 871,025 specimens Vision
Virchow [23] Self-Distillation DINOv2 ViT-H Memorial Sloan Kettering Cancer Center (MSKCC) Slides: 1,488,550 Tiles: 2 billion H&E 17 organs 119,629 patients 208,815 cases 392,268 specimens Vision
RudolfV [64] Self-Distillation DINOv2 ViT-L 108,433 slides from 15 labs across the EU and US + 26,565 slides from TCGA Slides: 133,998 Tiles: 1.2 billion H&E (68%) + IHC (15%) + other (17%) 14 organs 15 lab 58 tissue types 129 stains 6 scanner types FFPE and FF tissue Vision
PLUTO [65] Self-Distillation+MIM DINOv2 variation + MAE ViT-S variant FlexiViT-S [87] TCGA + Proprietary data from PathAI Slides: 195 million Tiles: 158,000 H&E + IHC + Other stains More than 30 diseases More than 12 scanners More than 100 stain types More than 4M pathologist pixel-level annotation Vision
Hibou [88] Self-Distillation DINOv2 ViT-B (for Hibou-B) ViT-L (for Hibou-L) Proprietary Data Slides: 936,441 H&E +202,464 non-H&E + 2,676 cytology slides Tiles: 1.2 billion patches for Hibou-L, 512 million patches for Hibou-B H&E + non H&E 306,400 cases Vision
BEPH [79] MIM BEiTv2 [89] with VQ-KD autoencoder and ViT-B encoder TCGA Slides: 11,760 Tiles: 11,774,353 H&E 32 cancer types Vision
UNI [90] Self-Distillation DINOv2 (also MoCov3 for comparison) ViT-L Proprietary Data from Massachusetts General Hospital (MGH), Brigham and Women’s Hospital (BWH), Genotype–Tissue Expression (GTEx) consortium Slides: 100,426 Tiles: over 100 million H&E 20 major tissue types Vision
3B-CPath [91] MIM and Self-Distillation MAE, DINO ViT-S, ViT-L Mount Sinai Health System (MSHS) Slides: 423,600 Tiles: 3.25 billion H&E 3 anatomic site 2 institutions Vision
PathoDuet [92] Contrastive Learning MoCov3 ViT-B TCGA Slides: 11,000 Tiles: 13,166,437 H&E Vision
Phikon [93] MIM+Self-Distillation iBOT ViT-S, ViT-B, ViT-L TCGA (Three variants TCGA-COAD, PanCancer4M, PanCancer40M) Slides: 6,093 Tiles: 43,374,634 (For PanCancer40M) H&E 16 cancer types 13 anatomic sites 5,558 patients (For PanCancer40M) Vision
CTransPath [94] Contrastive Learning MoCov3 variation (SRCL: semantically relevant contrastive learning ) Swin Transformer TCGA + PAIP Slides: 32,220 Tiles: 15.6 million H&E 32 cancer types 25 anatomic sites Vision
Prov-GigaPath [26] MIM and Self-Distillation MAE, DINOv2 GigaPath [26] (constructed with ViT and LongNet [95]) Providence Health System (PHS) Slides: 171,189 Tiles: 1.3 billion H&E IHC 30,000 patients 28 cancer centers 31 major tissue types Vision Language
CONCH [24] MIM + Self-Distillation iBOT ViT-B backbone in image encoder In-house dataset Slides: 21,442 Tiles: 1.3 billion Self-Distillaton 350 cancer subtypes Vision Language
The term “unimodal” is used to signify the pre-training phase utilizing a single modality out of vision and language. This is inherently different from vision-language pre-training strategies which involve both vision and language modalities to learn a joint vision-language representation space. In the unimodal vision pre-training phase there are three strategies commonly used in CPath which are self-distillation, contrastive learning and masked image modeling (MIM) approach. Each approach has its own advantages and disadvantages; however, the most popular approach is self-distillation with no labels (specifically DINOv2) which has been used in all the recent FMs [23, 64, 90, 88] as the SSL scheme. The work done in Lunit [96] goes through and benchmarks DINO, MoCov2, SwaV [97], Barlow Twins [98], but concludes there is no clear best SSL scheme. Lai [99] et al. is another work that performs a thorough evaluation of SSL schemes in CPath. .
Figure 10: Visualization of the self-distillation approach of unimodal vision pre-training scheme.
III-B1 Unimodal Vision Pre-training
TABLE V: Summary of Vision-Language Pre-Training and Instruction Tuning Phase of Foundation Models in Pathology
Model Vision Language Modules Vision-Language Pre-Training Phase Instruction Tuning Phase
Vision Module Language Module Additional Layer/Module Architecture / Framework Pre-Training Process Instruction Tuning Process
CONCH [24] An image encoder with ViT-B backbone with 12 transformer layers, 12 attention heads followed by two attentional pooler modules A GPT-style text encoder with 12 transformer layers A GPT-style multimodal decoder with 12 transformer layers CoCa Visual-language pre-training with image–text contrastive loss and the captioning loss according to CoCa framework
PLIP [40] An image encoder with ViT-B A text transformer as the text encoder with modifications mentioned in [100] CLIP Fine-tuning of CLIP model through contrastive learning using OpenPath dataset
PathChat [25] UNI [90] as vision encoder backbone Pre-trained Llama 2 [101] LLM which is a decoder-only transformer-based auto-regressive model A multimodal projector module to connect the outputs of the vision module to language module by projecting the visual tokens to the same dimension as the LLM’s embedding space for text tokens CoCa Vision-language pre-training according to CoCa framework with CONCH dataset LLaVa 1.5 [63] training approach First phase: Only the parameters of multi-modal projector is updated Second phase: Fine-tuned with instruction-following data
PathCLIP and PathAsst [38] Image encoder with ViT-B Text transformer as the text encoder with modifications mentioned in [100] Fully-connected layer after vision encoder to map the image embedding space to the corresponding language embedding CLIP Fine-tuning of CLIP model through contrastive learning using PathCap dataset First phase: Detailed description-based part is used to train the fully-connected layer connected to the vision encoder. Second phase: fine-tuned with instruction-following data via next word prediction
PRISM [102] Virchow as tile encoder, Perceiver [103] as slide-encoder BioGPT first 12 layers BioGPT last 12 layers as vision-language decoder CoCa Trained using contrastive loss and generative/ captioning loss
[…]
my tier/.style=tier/.wrap pgfmath arg=level##1level(), , The self-distillation with no label approach uses a student-teacher network to learn a rich vision representation space as shown in Fig. 10. From a single image patch, two different views are generated by applying an augmentation sampled from a set of possible augmentations (color jittering, Gaussian blur, polarization, etc). The generated output of both the networks is utilized to compute a cross-entropy loss which is then used to update the parameters of the student network. The parameter of the teacher network is then updated through an exponential moving average (EMA) of the student network parameters. Among the surveyed articles Virchow2, Virchow, RudolfV, Hibou and UNI use DINOv2 as the self-distillation approach. PLUTO takes a unique approach by integrating MAE and a Fourier loss term to get a custom variation of DINOv2.
Figure 11: Visualization of the masked image modeling approach of unimodal vision pre-training scheme.
Figure 12: Visualization of the MoCo approach of unimodal vision pre-training scheme.
The second popular SSPT approach is the MIM (visualized in Fig. 11) which has variants like MAE and iBOT used in FMs. In the MAE approach, randomly selected high portions of the image are masked out and the patches that are not masked out are passed through an encoder which generates latent representations of those patches.
Then those representations are passed through a decoder along tokens of masked out regions to reconstruct the image and reconstruction loss is used to train the model. The iBOT approach also uses MIM but adds self-distillation technique by leveraging a student-teacher network. The teacher network works as an online tokenizer and the student network learns to predict masked patches with the help of distilled knowledge of the teacher network.
Among the surveyed articles 3B-CPath and Prov-GigaPath use MAE but in conjunction with DINO and DINOv2 respectively. Research works utilizing iBOT approach include Phikon and CONCH.
Another SSPT approach that is comparatively less popular in FMs for CPath is the contrastive learning framework proposed in MoCo [104] (visualization in Fig. 12). Over the years variations of the proposed approach in MoCo in terms of architectural change and training blueprint have been done and MoCov2 [105] and MoCov3 [83] are the results of that.
Like the self-distillation approach, MoCo also utilizes two models; one is an encoder (with query patch as input) and the second one is a special momentum encoder (with key patches as input). The embeddings generated through the encoder and the momentum encoder are used to compute a similarity score which in turn is used for contrastive loss computation. The computed loss is used in backpropagation to update the parameters of the encoder. The parameter of the momentum encoder is updated through a momentum-based update rule that utilizes the parameters of the encoder. Another innovation of MoCo was introducing a feature queue which ensures a large set of negative samples without holding the entire dataset in memory. This also ensures the negative samples do not go stale over the training period and the model sees a diverse set of negative samples to learn a better vision representation. Among the surveyed articles PathoDuet uses MoCov3, UNI uses MoCov3 for comparison and CTransPath performs a unique variation of MoCov3 called semantically relevant contrastive learning (SRCL).
GPFM [22] is distinct from other approaches as it includes a novel expert knowledge distillation in addition to MIM and self-distillation by utilizing existing FMs UNI, Phikon and CONCH.
III-B2 Vision-Language Pre-training
After unimodal pre-training for the image module and text module, VLFMs go through vision language SSPT to align or learn a joint vision-language representation space.
The most popular vision-language SSPT approach in CPath is the CLIP approach shown in Fig. 13.
The inputs to CLIP are a batch of image features encoded by an image encoder and a batch of text features encoded by a text encoder. The model is trained to pull together the corresponding image feature and text feature in the representation space and push apart the rest of the features i.e. utilizing a contrastive objective. In other words, maximizing the similarity (specifically cosine similarity) between the corresponding image and text pair while minimizing the cosine similarity with other non-matching embeddings in the representation space.
Figure 13: Visualization of the CLIP approach of the vision-language pre-training scheme.
The second widely used approach is the CoCa [86] framework shown in Fig. 14.
Figure 14: Visualization of the CoCa approach of the vision-language pre-training scheme.
It contains three different modules; an image encoder, an unimodal text encoder and a multi-modal text decoder. It is trained by leveraging both captioning loss which follows a generative objective and contrastive loss which follows a contrastive objective. The input image patches are passed through an image encoder to generate image embeddings which is used to provide cross-attention to the multi-modal text decoder. The combination of generative and contrastive objectives forces the model to learn rich vision-language representation space.
Details of vision modules, language modules, additional modules and specifics of the pre-training process of FMs are outlined in Table V.
One recent work Zhou et al. [^1] https://github.com/MAGIC-AI4Med/KEP [106] proposes an entirely new approach in visual-language SSPT named knowledge enhanced pre-training (KEP) by utilizing PathKT, a pathology knowledge tree curated by the same research. This is different from the CLIP and CoCa pre-training techniques as it uses a novel knowledge encoder and knowledge distillation technique.
III-C Instruction-Tuning Phase
For CPath, the instruction-tuning phase provides the model with conversational ability i.e. a user can prompt the model and the model will respond according to the prompt. Visual-instruction tuning adds the ability to provide an image in addition to natural language user prompts. Among the surveyed papers in FMs, PathChat and PathAsst perform instruction-tuning with self-curated datasets PathChatInstruct and PathInstruct, respectively. Both works adopt the strategy employed by LLaVA and LLaVa-1.5 which are the pioneering work in visual instruction tuning. A summary of the instruction tuning process is provided in Table V. For both PathChat and PathAsst, the instruction tuning phase is subdivided into two phases. For both models in the first phase, the parameters of the vision encoder are frozen and the layer/module connected to the vision encoder (for PathChat it is a multimodal projector, for PathAsst it is fully connected layers) is trained. In the second phase, the model is fine-tuned with instruction-following data.
Even though instruction tuning in CPath is relatively new, some recent works like PathInsight [107] specifically focus on this aspect of FMs. Another recent work CLOVER [^2] https://github.com/JLINEkai/CLOVER [108] provides a framework for cost-effective instruction learning in pathology.
III-D Downstream Tasks
FMs have the ability to adapt to a vast array of tasks by utilizing the representation space learned during SSPT.
Figure 15: Adaptation and performance evaluation strategy for downstream tasks.
Note that, in the pre-training phase FMs were never trained for any of these specific tasks. At the end of the pre-training of FMs, one of the following strategies is adopted to perform a specific task.
  1. 1.
    Linear Probing: This is a commonly used technique where a linear classifier/regressor is trained on top of the pre-trained model. During the training of the linear layers, the parameters of the pre-trained model are kept frozen. Depending on the specifics of the tasks, the corresponding loss function and update rule are determined. This is a computationally cheap way to adapt to a downstream task as the parameters of the pre-trained model do not need to be updated.
  2. 2.
    KNN Probing: This is yet another approach to adopt a foundation model for a specific downstream task by utilizing K-nearest neighbors algorithm.
  3. 3.
    Fine-Tuning: This is similar to linear probing as a classifier/regressor is added on top of the pre-trained model, but the major difference is the parameter of the pre-trained models is also updated during fine-tuning. Hence, it is computationally much more costly compared to linear probing. This is sometimes also referred to as the supervised training phase. Note that, most of the time annotation is only available on slide-level. Hence, if the slide-level label is utilized it is called weakly supervised training.
TABLE VI: Different types of downstream tasks and corresponding research works
Downstream tasks Performed By
Cancer Detection [40],[102],[92],[91],[79],[23]
Tumor Detection [40], [38],[90]
Disease/Cancer/Tissue/Tumor/Molecular SubTyping [40], [24],[102],[26],[92],[90],[79],[64]
Cancer Grading [24],[90]
Image/Tissue/Tumor/Nuclear Segmentation [24],[90],[64]
Survival Prediction [79]
Text-to-Image Retrieval [40],[24]
Image-to-Text Retrieval [24]
Image-to-Image Retrieval [40],[90],[64],[22]
Image Captioning [24]
Pattern/Tissue/Image Classification [24],[38],[90]
Biomarker Prediction/Detection/Screening/Scoring [102],[90],[91],[64],[23]
Metastasis Detection [90]
Organ Transplant Assessment [90]
Mutation Detection/Prediction [26],[90],[91]
VQA [38],[25],[22]
Report Generation [102],[22]
Survival Analysis [22]
Conversational Agent [38],[25]
In Table VI, a summary of downstream tasks is provided along with research works performing these tasks.
Figure 16: Aggregate patch-level features through an aggregator network to get a slide-level representation and use it for weakly supervised training.
Another aspect to consider is how the performance evaluation is conducted. Among the surveyed articles, there are three different strategies that are employed.
  1. 1.
    Zero Shot Evaluation: In zero shot evaluation, the pre-trained model is directly used in a downstream task without probing or fine-tuning the pre-trained model with any samples of the downstream dataset. This provides a direct assessment of learned representation in the pre-training phase i.e. evaluates the quality of the generated embeddings from the pre-trained model. This is the most common approach among the surveyed articles.
  2. 2.
    Few Shot Evaluation: In few shot evaluation, the pre-trained model sees only a few examples from the downstream task dataset.
  3. 3.
    Simple Shot Evaluation: This is a unique variation of the few shot evaluation methods which is only been explored in UNI.
III-E Framework, Benchmarking and Adaptation of FMs
There are several research works in CPath that do not directly propose a FM but introduce the framework of FMs, provide benchmarks and comparisons between FMs and adapt the FMs for efficient training.
TABLE VII: Evaluation datasets and the associated models utilizing the dataset along with the data source
Dataset Used By Source
PanNuke [109] [40]
DigestPath [110] [40], [24]
WSSS4LUAD [111] [40],[38],[24]
SkinCancer [112] [39]
PatchCamelyon (PCam) [113] [88],[64],[23],[39]
MHIST [114] [88],[64],[23],[39]
BACH [115] [79],[90],[39]
SICAP, SICAPv2 [116] [24],[39]
Databiox [117] [39]
RenalCell [118] [39]
Osteo [119] [39]
CoNSeP [120] [23]
KIMIA Path24C [121] [40]
DHMC [122] [24],[90]
EBRAINS [123] [24],[90]
AGGC [124] [24],[90]
PANDA [125] [24],[90]
MSK-IMPACT [126] [102],[23]
SegPath [127] [90]
BRACS [128] [90]
UniToPatho [129] [90]
HunCRC [130] [90]
BreakHis [131] [79]
MSI-CRC & MSI-STAD [132, 133] [88],[64]
TIL-DET [134] [88],[64]
MIDOG [135] [23]
CAMELYON16 [136] [79],[92],[90]
CAMELYON17-WILDS [137] [90],[23]
[…]
[38],[24],[90],[88],[64]
[79],[23], [39]
[23]
[…]
[38],[39]
[38],[39]
TCGA Uniform [140] [90]
TCGA CRC-MSI [90],[23]
TCGA-TILs [141] [90],[23]
[…]
[24],[102],[79],[88]
[24],[79],[88],[92]
[24],[102],[79],[88],[92]
Mentioned data sources are shared through platforms like Zenodo, GitHub, official challenge websites. TCGA dataset needs to be downloaded from the official portal.
III-E1 Frameworks:
eva [^3] https://github.com/kaiko-ai/eva [71, 142] is a framework for VFMs in CPath which abstracts a lot of complexity of VFMs. In addition, it facilitates the reproducibility of VFMs for fair comparison and provides an interface to evaluate publicly available downstream datasets. Note that, this is an ongoing work and more downstream datasets and models are being added.
III-E2 Benchmarking:
Another category is benchmark analysis of FMs [67], [66],[143] in CPath. The work done in [67] analyzes different FMs and evaluates their performance on $88$ datasets. The other work [66] is a more recent work that analyzes the performance of FMs on a large and diverse data set collected from two medical centers. The FMs were benchmarked on $33$ broad downstream tasks which include disease detection, biomarker prediction and treatment outcome prediction.
III-E3 Adaptation of FMs:
Another category is the adaptation of existing FMs to carry out tasks such as low-resource fine-tuning[72] and multi-modal prompt-tuning[73]. The work carried out in [72] fine-tunes a FM with a single GPU and shows that it can outperform SOTA feature extractors. PathoTune [73] adapts a visual or pathological FM to downstream tasks using multi-modal prompts.
Other than the surveyed FMs, there are more recent works of FM in CPath like mSTAR [144]. However, as it includes RNA-Seq data in addition to pathology reports it falls outside the scope of this paper. Another foundation model we do not include is H-optimus-0444https://huggingface.co/bioptimus/H-optimus-0 [145] developed by Bioptimus as it is not associated with a publication.
TABLE VIII: Summary of Vision-Language Models in Computational Pathology
Model Architecture and Utilized Models/Frameworks Dataset Contribution Availability
TraP-VQA [146] Question feature extraction with BioELMo [147] and BiLSTM Image feature extraction with ResNet-50 and CNN layers Transformer encoder to fuse question and image feature, transformer decoder to upsample fused features PathVQA Performs VQA Provides interpretability in text domain with SHAP [148] Provides interpretability in image domain with GradCAM [149]
FSWC [150] CLIP for image and text feature extraction GPT-4 to generate instance-level and slide-level prompt groups which introduce pathological prior knowledge into the model Camelyon 16 TCGA-Lung In-house cervical cancer dataset Performs few-shot weakly supervised WSI classification Proposes a two-level (instance-level and slide-level) prompt learning MIL framework named TOP Introduces a prompt guided instance pooling to generate slide-level feature
ViLA-MIL [151] GPT-3.5 as the frozen LLM which generates visual descriptive text for WSI at 5x and 10x resolution based on class level question prompt ResNet-50 as the image encoder and corresponding CLIP transformer is used as the text encoder Prototype-guided patch decoder which generates slide features and context-guided text decoder which generates text-features TIHD-RCC (in-house dataset) TCGA-RCC TCGA-Lung Performs WSI classification with a MIL framework that utilizes information from 5x and 10x. In addition, incorporates information from LLM generated visual descriptive text for both 5x and 10x scale WSIs Introduces a novel prototype-guided patch decoder that progressively aggregates the patch features Introduces a context-guided text decoder to refine the text prompt features by leveraging multi-granular image contexts
PathM3 [152] ViT-G as the image encoder Query-based transformer to fuse image embeddings of WSIs and corresponding captions Frozen LLM FlanT5 [153] for caption generation PatchGastric [154] Performs WSI classification and captioning through a multi-modal, multi-task, multi-instance learning framework Leverages limited WSI captions during training Develops multi-task joint learning inspired from [155]
HLSS [156] ResNet-50 as visual encoder with CLIP pre-training Language encoder from CLIP A positive pairing module (PPM) which consists of three parallel reshape layer followed by MLP A cross-modal alignment (CAM) module which computes cosine similarity OpenSRH [157] TCGA Performs hierarchical (patient-slide-patch hierarchy) text-to-vision alignment for SSL Proposes a positive pairing module (PPM) and a cross-modal alignment module (CAM)
FiVE [158] ResNet as image encoder following [159] Pre-trained BioClinicalBERT [160] as text encoder and utilizing LoRA [161] for fine-tuning GPT-4 to generate fine-grained pathological descriptions based on raw pathology reports and expert inquiries An instance aggregator module consisting of a self-attention module and a cross-attention module that fuses image instance embeddings and prompt embeddings to create bag-level feature Camelyon 16 TCGA-Lung Incorporates a patch-samnpling strategy for optimize training efficiency of the model Proposes a task-specific fine-grained semantics (TFS) module Performs zero shot histological subtype classification, few shot classification and supervised classification with pre-training
CPLIP [162] MI-Zero [62] to identity-related prompts based on an image by utilizing similarity metric GPT-3 to categorize and transform prompts into five variations PILP to match the transformed prompts with corresponding images from OpenPath CRC100K WSSS4LUAD PanNuke DigestPath SICAP Camelyon 16 TCGA-BRCA TCGA-RCC TCGA-NSCLC Creates a pathology-specific dictionary using a range of publicly available online glossaries Introduces a many-to-many contrastive learning instead of traditional one-to-one contrastive learning
HistGen [163] GPT-4 to clean and summarize reports curated from TCGA A ViT-L model is pre-trained with 200 million patches by utilizing DINOv2 and used as feature extractor UBC-OCEAN [164, 165] TUPAC16 [166] Camelyon 16 and 17 TCGA-BRCA TCGA-STAD TCGA-KIRC TCGA-KIRP TCGA-LUAD TCGA-COADREAD Performs WSI report generation, cancer subtyping and survival analysis Develops a local-global hierarchical encoder module to capture features at different levels (region-to-slide) Develops a cross-modal context-aware learning module to align and ensure interaction between different modalities
CLIPath [167] ResNet-50 as the vision encoder in CLIP $44$ fully-connected layer after the vision encoder Transformer as the text encoder in CLIP PatchCamelyon (PCam) MHIST Develops residual feature connection (RFC) to fine-tune CLIP with a small amount of trainable parameters and this also fuses the existing knowledge from pre-trained CLIP and the newly learned knowledge related to pathological tasks
MI-Zero [62] HistPathGPT: Unimodal pre-training of GPT-style transformer (same architecture as GPT 2-medium [100]) which is used as text encoder. Additionally, BioClinicalBERT [168] and PubMedBERT [169] are considered as text encoders. Existing encoders like CTransPath [94] (which is based on Swin Transformer [170]) or ViT-S (ImageNet initialization or pre-trained with MoCov3) 3 WSI datasets from Brigham and Women’s Hospital (in-house dataset) named Independent BRCA, Independent NSCLC, Independent RCC TCGA-BRCA TCGA-NSCLC TCGA-RCC Performs cancer subtyping with a VLM Develops a custom pathology domain-specific pre-trained text encoder called HistPathGPT Introduces 33,480 image-caption pair dataset
HistoGPT [171] The vision module is based on CTransPath BioGPT as the language module Image features sampled (with Perceiver Resampler [103]) from the vision modules are integrated into the language module via interleaved gated cross-attention (XATTN) blocks [172] In-house dataset with 2 cohorts. Munich cohort with l 6,000 histology samples and Münster cohort with 1,300 histology samples with all samples stained with H&E Performs histopathology report generation that provides a description of WSIs with high fidelity Provides interpretability map which shows which word in the generated report corresponds to which region in a WSI
Model Architecture and Utilized Models/Frameworks Dataset Contribution Availability
PathAlign [173] PathSSL patch encoder which is based on ViT-S architecture and pre-trained following the approach in [99] with masked siamese networks [174] SSL scheme Q-Former from BLIP-2 [175] as WSI-encoder Among two variants PathAlign-R and PathAlign-G, PathAlign-G uses PaLM-2 S [176] as the frozen LLM In-house de-identified dataset (DS1) of WSIs paired with reports collected from a teaching hospital. The stain type for the WSIs are H&E and IHC TCGA Performs WSI classification, image-to-text retrieval and text generation
MI-Gen [34] ResNet and ViT (ImageNet initialization or hierarchical SSPT with HIPT [177]) as visual encoder CNN layer in the hierarchical position-aware module (PAM) Transformer as encoder-decoder (both vanilla transformer and Mem-Transformer [178]). Additionally, CNN-RNN [179] and att-LSTM [180] is used for comparison PathText dataset which was created using and WSI and pathology report from TCGA-BRCA Performs pathology report generation from WSI Additionally performs cancer subtyping and biomarker prediction Develops a hierarchical position-aware module (PAM) Introduces PathText dataset which contains 9,009 WSI-text pairs
W2T [35] ResNet-50, ViT-S pre-trained with DINO or scheme followed in HIPT as visual encoders to extract patch features Text Encoders: PubMedBERT and BioClinicalBERT Co-attention mapping in the decoder to align visual and text features WSI-VQA Curates a WSI VQA dataset called WSI-VQA with 977 WSIs and 8,672 Q/A Performs VQA with the proposed WSI-VQA dataset Co-attention mapping between word embeddings and WSIs are used for interpretability
PathGen-CLIP [37] PathGen-LLaVA OpenCLIP framework is utilized to train a pathology-specific CLIP model The vision encoder of LLaVA v1.5 is replaced with the trained pathology-specific CLIP model PathGen-CLIP is combined with Vicuna LLM PathGen-1.6M (for pre-training) PatchCamelyon (Pcam) CRC-100K SICAPv2 BACH Osteo SkinCancer LC25000 Camelyon 16 and 17 Curates a large-scale dataset PathGen-1.6M with 1.6 million image-caption pairs Develops a pathology-specific large multi-modal modal with the capability to adapt to various downstream task
Quilt-Net [39] CLIP model (based on OpenCLIP framework) ViT-B as the image encoder GPT-2 and PubMedBERT as text encoder Quilt-1M (for pre-training) PatchCamelyon (PCam) NCT-CRC-HE-100K SICAPv2 Databiox BACH Osteo Renal Cell SkinCancer MHIST LC25000 Curates a large-scale dataset Quilt-1M Performs zero-shot, linear-probing task and also cross-modal retrieval task Provides cross-modal attention mask with Grad-CAM
Guevara et al. [181] CLIP model HIPT as image encoder A transformer encoder to project encoded image embedding to match the dimension of text embedding PubMedBERT as text encoder For the caption model pre-trained Bio GPT 2 is used GPT-3.5-turbo [182] to clean and refine captions. Additionally, the captions were machine-translated from two languages to English using the approach in [183] Self-curated dataset of WSIs and captions. WSIs are of colon polyps and biopsies. The captions contain possible 5 diagnostic labels normal, hyperplasia, low-grade dysplasia, high-grade dysplasia, or adenocarcinoma Performs WSI classification and caption generation through the use of weakly supervised transformer-based models
Hu et al. [184] ResNet-50 pre-trained with BYOL [185] SSL scheme as patch encoder Anchor-based module as WSI encoder based on [186] which uses kernel attention transformer Prompt-based text encoder which utilizes self-attention structure in BERT [187] GastricADC [188] In-house gastric dataset containing 3598 WSIs and reports dataset Performs cross-modal retrieval tasks which includes image-to-image retrieval, image-to-text retrieval, text-to-image retrieval and text-to-text retrieval Proposes a histopathology language-image representation learning framework Develops a prompt-based text representation learning scheme
PEMP [189] CLIP as the backbone of the proposed model A self-attention layer A attention pooling layer In-house datasets TCGA-CESC Performs survival analysis, metastasis detection and cancer subtype classification Proposes a mechanism to introduce vision and text prior knowledge in the designed prompts (both static and learnable) at both patch and slide levels Develops a self-attention layer called a lightweight messenger and an attention pooling layer called summary layer
HistoCap [190] and HistoCapBERT ResNet-18 to encode thumbnail image LSTM decoder according to [180] as caption generator Pre-trained HIPT encoder for WSI encoding Trainable attention layer after HIPT encoder In HistocapBERT the LSTM decoder is replaced by a BioclinicalBERT Dataset from Genotype-Tissue Expression (GTEx) portal Performs caption generation given a thumbnail image and a WSI image
PathCap [191] ResNet-18 to encode thumbnail image ResNet-18 to encode WSI patches Trainable attention layer after the ResNet-18 used to encode WSI patches LSTM as caption generation module Dataset from Genotype-Tissue Ex- pression (GTEx) portal Performs caption generation of histopathology images using multi-scale view (thumbnail of WSIs and WSIs)
Elbedwehy et al. [192] Vision encoders include VGG [193], ResNet, PVT [194], Swin-Large [170], ConvNexT-Large [52] Language decoders and pre-trained word embedding models include LSTM, RNN, Bi-directional RNN, BioLinkBERT-Large [195] PatchGastric Performs caption generation of WSIs
PromptBio [196] PLIP as coarse-grained pathology instance classifier which filters out patches given a patch and a prompt (keeps patches associated with cancer-associated stroma) IBMIL [197] to encode patches followed by a fully-connected layer GPT-4 to generate pathology descriptions Series of transformer and MLP layer which performs biomarker prediction TCGA-CRC The Clinical Proteomic Tumor Analysis Consortium (CPTAC) CRC dataset Performs biomarker prediction
Model Architecture and Utilized Models/Frameworks Dataset Contribution Availability
SGMT [198] ResNet-18 (pre-trained on ImageNet) as patch encoder Transformer encoder to encode the patch features Transformer decoder as caption generation module PatchGastric Performs caption generation of WSIs Proposes a novel mechanism to use subtype prediction as a guiding mechanism for the caption generation task Develops a random sampling and voting strategy to select patches
Tsuneki et al. [154] EfficientNetB3 [199] and DenseNet121 [200] (pre-trained on ImageNet) as image encoder Average and global pooling layer after the image encoders RNN decoder layer for caption generation PatchGastric Performs caption generation of WSIs Curates a captioned dataset of 262,777 patches from 991 WSIs
IV Vision-Language Models
In this section, the recent works in CPath with VLMs are outlined. The details about architecture, used datasets and contribution of individual research work are listed in Table VIII.
First, different categorizations of VLMs are provided to give insight into how VLMs are utilized to solve pathology-specific tasks. Second, the common architectural components in VLMs and adopted strategies are summarized. Lastly, a brief overview of models that do not solve a direct pathology task but perform other types of vision-language tasks is summarized.
IV-A Categorization of VLMs
A categorization of VLMs in CPath can be done based on the reason for using the language modality. Some research works solve tasks like caption generation ([191, 154, 198, 190, 192]) or VQA ([146, 35]) which necessitates utilization of both vision and language modality because of the nature of tasks. For VQA or caption generation, the generated output from the model needs to be in language form i.e. the model needs to perform a language task. On the other hand, other research works use language modality as a source of semantic information to be injected into information gained from vision modality. This additional semantic information can significantly boost the performance of the model which would not have been possible with a vision-only model. As an example, MI-Zero [62] performs cancer subtyping which is not a language task but utilizes pathology reports curated from the in-house dataset and TCGA as a source of semantic information.
Figure 17: Visualization of categorization of VLMs and VLFMs. There are no clear distinctive criteria for a model to be classified as a VLFM or a VLM.
Another categorization can be done by comparing the training approach with FMs. VLFMs are different from VLMs as traditional VLMs focus on solving one or two vision-language tasks like caption generation. However, as FMs become more prevalent recent works are starting to shift towards VLFMs. Recent works (PathGen-CLIP [37], Quilt-Net [39]) take FM-like approach to train their models and adapt them to different downstream tasks. On the other end of the spectrum, there are VLMs that solve only a single task. As shown in Fig. 17 VLMs and VLFMs should be viewed as a spectrum where if the number of tasks and pre-training data size increases it moves away from VLMs and towards VLFMs.
TABLE IX: Summary of pre-training strategies in VLMs
Strategy Reference
Performs pre-training from scratch (vision or language module) [39],[37]
Utilize domain-specific vision module (HIPT, CTransPath, etc) or language module (PubMedBERT, BioClinicalBERT, etc) [196],[190],[184], [181],[35], [34], [171]
Initialize with out-of-domain encoders (ImageNet pre-training, pre-trained CLIP, etc) [154], [198], [192] [156]
Owing to this, some VLMs follow the FM approach of pre-training the vision or text module. However, as pre-training from scratch requires a large amount of data and computing some VLMs use domain-specific vision modules from HIPT, CTransPath, IBMIL, etc or language modules from PubMedBERT, BioClinicalBERT, etc. This allows the vision or language module to learn an initial vision or language representation space which might not be as rich as pre-training scratch but boosts the performance.
Figure 18: Amount of semantic information available through different text sources.
Another categorization can be done by analyzing how language prior is incorporated into the model. Most works use text labels or captions curated from different sources, which provide the least amount of information. Recent works like FSWC [150], ViLA-MIL [151] and PEMP [189] utilize prompting LLMs to increase the amount of semantic information that can be gained from text labels or captions. They prompt the LLMs to generate descriptions about a particular class label or of morphological or textural patterns about patches. On the other end of the spectrum is research work utilizing all information from pathology reports which contain a lot of details. However, as shown in Fig. 18 the difficulty of curation and doing it at a scale to match the pathologist level also increases.
IV-B Architectural components for VLMs in CPath
Even though there is a huge variety of vision and language modules used in VLMs, it is possible to identify common structures. In Fig. 19 the common architectural components used in different stages are outlined. In the pre-processing stage, LLMs like GPT-4, GPT-3.5, GPT-3.5-turbo and GPT-3 are used to clean and refine captions or prompting them to generate descriptions. Examples of research work utilizing LLMs in the pre-processing phase include HistGen, FSWC, ViLA-MIL, FiVE, CPLIP and PromptBio.
Another key component of the architecture is the vision module or encoder which converts the WSI patches to image embeddings. As shown in Fig. 19 it can be separated into three separate groups. The first group is vanilla CNNs that have ImageNet initialization or some kind of pre-training on pathology data. Note that most VLMs utilizing these architectures are not recent works (with the exception of Elbedwehy et al. [192]). Recent VLMs utilize ViTs which is the second group containing several ViT variants. The third group is special encoders that are pathology domain-specific and optimized for encoding WSI patches. These are ideal for learning a rich representation without the cost of pre-training from scratch.
For language encoders, most works utilize BERT (PubMedBERT, BioClincialBERT) or GPT (BioGPT, GPT 2) variants and most are pre-trained on datasets from the biomedical domain.
Another type of language module is the caption/text generation module which rather than encoding text into embedding, generates text sequences. Earlier works included RNN, LSTM and Transformer decoders and recent works include LLMs.
The most common vision-language alignment or fusion module is the CLIP model and its variations. However, some works come up with custom approaches to align vision and language representations.
Figure 19: Common modules used in different stages of VLMs. The last stage is the vision-language alignment or fusion stage where the vision and language embeddings generated from the earlier stages are combined.
IV-C Related works for VLMs in CPath
There are several research works that do not directly utilize VLMs to solve a task in the pathology domain, but investigate or use VLMs for other reasons. One example is Thota et al  [^5] https://github.com/jaiprakash1824/VLM_Adv_Attack [201] which investigates the effect of Projected Gradient Descent (PGD) adversarial perturbation attack on PLIP [40] architecture. Another work by Lucassen et al. [^6] https://github.com/RTLucassen/report_preprocessing [202] provides a pathology report processing workflow that can be used for VLMs in CPath.
Another category of VLMs in CPath that does not directly solve pathology-specific tasks is text-guided diffusion models for image generation. However, these models can perform tasks like virtual stain transfer which can later be used to solve a pathology-specific task. Another use case is extending dataset size with generated synthetic images. One such work is PathLDM [^7] https://github.com/cvlab-stonybrook/PathLDM [203], that performs text-to-image generation on the TCGA-BRCA dataset.
V Conclusion
In recent years the number of research works in CPath with FMs and VLMs has increased significantly which provides an indication about how CPath will evolve in the next couple of years. Many of these works put significant computing resources into training FMs on massive pre-training datasets or coming up with novel strategies to push more language prior knowledge into VLMs. In the near future, the VLFMs which bring together the benefits of both FMs and VLMs, are going to be the dominant model. This review article provides a comprehensive overview of all these models which will aid future researchers.